Find the cost per ounce of a mixture of 240 oz of a cologne that costs $6.50 per ounce and 560 oz of a cologne that costs $4.00 per ounce.
Answer by fractalier(6550)

Call the cost you wish to find, x. The set up is like this
240(6.50) + 560(4.00) = (240+560)x
1560 + 2240 = 800x
3800 = 800x
x = $4.75 per ounce